Fermi National Laboratory
 

Operation Summary for February 8, 2010 - February 10, 2010

 

Monday Feb. 8

The day shift began with the Tevatron (TeV) in access, with Operations monitoring the Recycler stash (336.0E10), with the Antiproton Source stacking (12.0E10), with the Linac, Booster, and the Main Injector (MI) operating normally, but with no beam for the NuMI, MiniBooNE, and Meson MTest experiments, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, Operations reported numerous MI vacuum pump (I:IP606D) trips that caused occasional vacuum cascades throughout the day shift. TeV personnel accessed the TeV at the end of the last shift to diagnose a bad lead. They completed their work and Operations began the TeV turn-on at 10:15 AM. Operations put the TeV into HEP shot setup at 3:03 PM.

During the evening shift, Operations established store 7589 at 4:15 PM with an Average Initial Luminosity (AIL) of 264.5E30. Operations established beam to MTest around 5:30 PM and the experiment took beam until 6:20 PM.

Tuesday Feb. 9

The midnight shift began with Operations monitoring store 7589 (102.67E30) and the stash (325.4E10), with Pbar stacking (16.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7589 at 1:03 AM and then put the TeV into shot setup. Operations established store 7590 at 2:51 AM with an AIL of 257.7E30. Operations had MI, NuMI, and MiniBooNE turned off by 6:30 AM for scheduled accesses.

During the day shift, Operations reported that the Feynman Computer Center experienced a power outage at 12:45 PM. The MI RF cavity change-out and other jobs ran throughout the shift.


Feynman Computer Center

During the evening shift, RF personnel completed the MI cavity work and Operations turned MI back on at 10:43 PM. Operations reestablished beam to Pbar, NuMI, and MiniBooNE.

Wednesday Feb. 10

The midnight shift began with Operations monitoring store 7590 (45.32E30) and the stash (81.8E10), with Pbar stacking (7.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TeV quenched at TeV sector E1 and store 7590 was lost due to an earthquake centered near Virgil, IL. Operations prepared the TeV for scheduled safety system testing and accesses, which started at 7:30 AM.

The following plot shows various readouts during the earthquake.

The Plans for Wednesday

The plans for today are to complete the safety system testing in the TeV and then continue with normal stacking, stashing, and storing. The Run Coordinator will allow CDF to access for maintenance work and for Pbar and MI personnel to conduct studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/10/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory